calculate the density of carbon dioxide in grams per liter (g/L) at 0.990 atm and 55C

Gas Density Example #1 Calculate the density of carbon dioxide in grams per liter (g/L) at 0.990 atm and 55°C.
Answer #1

P = 0.99atm

T= 55.0 oC

= (55.0+273) K

= 328 K

Molar mass of CO2,

MM = 1*MM(C) + 2*MM(O)

= 1*12.01 + 2*16.0

= 44.01 g/mol

Lets derive the equation to be used

use:

p*V=n*R*T

p*V=(mass/molar mass)*R*T

p*molar mass=(mass/V)*R*T

p*molar mass=density*R*T

Put Values:

0.99 atm *44.01 g/mol = density * 0.08206 atm.L/mol.K *328.0 K

density = 1.62 g/L

Answer: 1.62 g/L